If stronium-90 has a half-life of fourteen days, and you have a sample that contains 30 kg of strontium-90, how much would remain after thirty-five days?
Answer by ankor@dixie-net.com(22740)

If stronium-90 has a half-life of fourteen days, and you have a sample that contains 30 kg of strontium-90, how much would remain after thirty-five days?
:
A = Ao*2^(-t/h)
:
A = 30*2^(-35/14)
A = 30*2^-2.5
A = 30*.17678
A = 5.3 kg after 35 days